Sol Summary – December 6th

Crew 184 Daily Summary Report  06 Dec 2017

MDRS Daily Summary Report for Sol 5

Summary Title:  Continuing failures, schedule replans, and a successful EVA!

Mission Status:  We are still waiting for a nominal morning!  To fix yesterday’s bandwidth problems we powered off all devices in the HAB, we then connected our ‘internet’ laptop to the modem and watched 100MB of bandwidth disappear in ~ 5 minutes.  We concluded there was a background process of some sort operating on that computer as no known programs were running.  We were able to fix this issue by swapping out this laptop and powering down the fault one.  Communication restored!

We exulted in this victory for approximately 10 minutes, then noticed that our generator had failed again and we were back to a powerdown configuration.  Our SOC was holding at 47%, but it was again clear that we were operating with a degraded power system.  Communication with Mission Control has tentatively identified the problem as due to a ‘summer fuel’ blend which has been noted before, which is separate from the oil leak we observed earlier.  The generator will likely continue to be faulty for the next few days until we can procure a ‘winter’ blend.  Winters on Mars can be tough!.  We are continuing to conserve power wherever possible and using the generator minimally to supplement the solar arrays when required.

We resumed nominal operations this afternoon with two primary activities

We decided to finish up the EVA that we left incomplete yesterday, going to Matryoshka Site #2 which is adjacent to the MDRS HAB.  Our EVA prep procedure continues to be challenging, but we are getting better.  We have allocated the best radios and headsets for EVA (recently delivered!) and placed them adjacent to the EVA prep room.  We have also pre-staged all EVA equipment (boots / gloves / first aid kid / emergency water supply / maps / etc . . . ) in the EVA room to minimize the chance that our prebreathe is interrupted by having to go out to the HAB to get supplies.  We have also started carrying backup radios with us since we have had to cut an EVA short due to radios running out of batteries.  The EVA training continues!

Once outside we successfully found samples from numerous geological strata in our target era and bagged them for later delivery and chemical analysis.  We had some excess time upon returning so we also took some of the PR images requested and will uplink them as part of the journalist report.

When the majority of the crew got back to the habitat we did an observatory training sessions where the crew astronomy gave a tutorial in observatory and telescope operation.  There’s something about being on Mars that gives observatory operations extra gravitas as you are already in the ‘space’ mindset.

Sol Activity Summary:

1.      Matryoshka EVA Site #2

2.      Observatory Training Session (no observing)

3.      HAB Cleaning

4.      Reports and general HAB maintenance / checks

5.      Food inventory transfer to excel

6.      Exercise

Look Ahead Plan:

  1. Matryoshka EVA Site #3
  2. Nerve Block Training
  3. CBS News Visit (unconfirmed – request confirmation on whether media crew is coming from Mission Support)
  4. Filming and Interviews
  5. Solar Observation
  6. Communication Experiment (Time Delay)
  7. CPR Technique Training
  8. Exercise / General HAB Maintenance

Anomalies in work:

1.      HAB Leak from under EVA door (per Mission Control leak is expected and has been present for months, presumed to be from Kitchen piping)

2.      Generator failure

3.      Internet bandwidth problem was resolved today.

Weather: Still cold (below freezing), slightly windy

Crew Physical Status: All crew in good physical health.

EVA: See above for EVA details

Reports to be filed:



EVA Request



Sol Summary

Support Requested:

1.      Trash pickup from rear airlock

2.      Replacement ATV Fuel

3.      Additional dish soap

4.  Confirmation of whether media visits will occur

Copyright © The Mars Society. All rights reserved. | Main Site